L-1{(e-2s)/[s2(s-1)]}


1
Expert's answer
2021-09-07T10:25:02-0400

"L^{-1}[\frac{e^{-2s}}{s^2(s-1)}]=L^{-1}[e^{-2s}(\frac{1}{s-1}-\frac{1}{s}-\frac{1}{s^2})]="

"=(e^t-1-t)\theta(t-2)."